# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ Abstract
Standard deep learning architectures used for classification generate label predictions with a projection head and softmax activation function. Although successful, these methods fail to leverage the relational information between samples in the batch for generating label predictions. In recent works, graph-based learning techniques, namely Laplace learning (Zhu et al. (2003)) or label propagation, have been heuristically combined with neural networks for both supervised and semi-supervised learning (SSL) tasks. However, prior works approximate the gradient of the loss function with respect to the graph learning algorithm or decouple the processes; end-to-end integration with neural networks is not achieved. In this work, we derive backpropagation equations, via the adjoint method, for inclusion of a general family of graph learning layers into a neural network. This allows us to precisely integrate graph Laplacian-based label propagation into a neural network layer, replacing a projection head and softmax activation function for classification tasks. Using this new framework, our experimental results demonstrate smooth label transitions across data, improved robustness to adversarial attacks, improved generalization, and improved training dynamics compared to the standard softmax-based approach [^1] Our code is available at https://github.com/jwcalder/GraphLearningLayer.
@@ 1 Introduction
Graph-based learning is a powerful class of transductive machine learning techniques that leverage the topology of a similarity graph constructed on the data for tasks such as semi-supervised learning, data visualization, and dimension reduction. Beginning with the seminal work by Zhu et al. (2003) on semi-supervised learning, graph-based learning techniques have since proliferated (Belkin and Niyogi, 2004; Belkin et al., 2004; Bengio et al., 2006; Zhou and Belkin, 2011; Wang et al., 2013; Ham et al., 2005; Lee et al., 2013; Calder et al., 2020; Calder and Slepcev, 2020; Shi et al., 2017) and found applications in many problem areas. Graphs may be intrinsically present in a data set, such as in network science problems, where relational information is present in links between servers in a network, or between users on social media platforms, or the data points themselves may be graphs (e.g., molecules in drug discovery). Another important graph representation is the similarity graph, which encodes the notion of similarity between pairs of data points. In this paper, we focus on graph-based learning using similarity graphs; that is, we do not assume the graph is given as part of the metadata in the problem. Thus, our work is orthogonal and complementary to much of the graph neural network literature, where the graph is a priori fixed.
@@
Similarity graphs and corresponding graph techniques have been extensively studied theoretically using variational analysis, random walks, harmonic functions, and partial differential equations (see e.g., Calder (2019); Calder and García Trillos (2022); Calder and Slepcev (2020); Calder (2018); Slepcev and Thorpe (2019); Dunlop et al. (2020); García Trillos et al. (2020); Hein et al. (2007, 2005); Calder and Ettehad (2022); Calder et al. (2023)). Graph learning also contrasts with modern deep learning approaches in that it is a data-dependent model without any trainable parameters. Previous works (Chapman et al., 2023; Chen et al., 2023a, b) show that graph learning outperforms many popular classifiers like support vector machine (Cortes and Vapnik (1995)), random forest (Ho (1995)), and neural networks on classification tasks, especially at very low label rates.
@@
Recently, many works have focused on combining graph-based learning with deep learning. In such works, unsupervised or semi-supervised neural networks learn an embedding of the data into a feature space and use the similarity between features to construction the graph. Examples of such deep learning techniques include variational autoencoders (VAEs) (Kingma and Welling (2013); Calder et al. (2020); Calder and Ettehad (2022)), convolutional neural networks (CNNs) (LeCun et al. (1989); Enwright et al. (2023)), convolutional VAEs (Pu et al. (2016); Miller et al. (2022); Chapman et al. (2023)), and contrastive learning (Chen et al. (2020a); Brown et al. (2023b, a)). Some works have used graph learning to generate pseudolabels to train a deep neural network in low label rate settings (Sellars et al. (2021); Iscen et al. (2019)). Another recent approach replaces the softmax activation with graph learning to improve generalization and adversarial robustness (Wang and Osher (2021)); however, the gradients of the loss with respect to the graph learning classifier are only approximated by an equivalently-sized linear classifier.
@@
A significant gap in the literature exists, as no work fully integrates graph learning into the gradient-based learning framework of neural networks. In this paper, we leverage the adjoint graph Laplace equation to derive the exact backpropagation gradients for a general family of graph learning methods, and show how to implement them efficiently. This significantly expands upon the previous work by Enwright et al. (2023), where the backpropagation equations were only derived for Laplace learning on a fully connected graph. The backpropagation equations allow us to design a neural network layer that replaces the standard projection head and softmax classifier with a graph learning classifier (see Figure 1). We call this novel layer the Graph Learning Layer (GLL). GLL can be combined with any neural network architecture and enables end-to-end training in a gradient-based optimization framework. Through a variety of experiments, we show that GLL learns the intrinsic geometry of data, improves the training dynamics and test accuracy of both shallow and deep networks, and is significantly more robust to adversarial attacks compared to a projection head and softmax classifier.
@@ 1.1 Related Work
Graph-based semi-supervised learning (SSL) methods construct a graph that allows inferences of unlabeled nodes using labeled nodes and the graph topology. The seminal work of Laplace learning (Zhu et al. (2003)), also known as label propagation, led to a proliferation of related techniques (Belkin and Niyogi, 2004; Belkin et al., 2004; Bengio et al., 2006; Zhou and Belkin, 2011; Wang et al., 2013; Ham et al., 2005; Lee et al., 2013; Garcia-Cardona et al., 2014; Calder and Slepcev, 2020; Shi et al., 2017). Laplace learning computes a graph harmonic function to extend the given labels to the remainder of the graph—essentially finding the smoothest function that agrees with the given labels. When the label rate is low (i.e., close to one label per class), Laplace learning degenerates and suffers from a “spiking” phenomenon around labeled points, while predictions are roughly constant (Calder et al. (2020); El Alaoui et al. (2016); Nadler et al. (2009)). Several methods have been proposed for these very low label rate regimes including reweighted Laplace learning (Shi et al. (2017); Calder and Slepcev (2020); Miller and Calder (2023)), $p$-Laplace methods (Zhou and Belkin (2011); El Alaoui et al. (2016); Flores et al. (2022); Slepcev and Thorpe (2019)), and Poisson learning (Calder et al. (2020)). Analysis of the degeneracy of Laplace learning from a random walk and variational perspective can be found in (Calder et al. (2020, 2023)). Some recent work has also focused on graph total variation problems using graph cuts (Bertozzi and Flenner (2016); Merkurjev et al. (2017, 2018)).
@@
Another related, but distinct, area of work in machine learning on graphs is graph neural networks, or GNNs (Kipf and Welling, 2016; Defferrard et al., 2016; Wu et al., 2020). The GNN literature considers how to incorporate a priori graph information relevant to the task, such as co-authorship in a citation graph, within a neural network architecture for various learning tasks, such as learning node embeddings or semi-supervised learning. These works are quite different. In particular, we do not assume the data has a graph structure. Moreover, GNNs contain trainable weight matrices, while graph-based learning methods have no trainable parameters. There are some works, such as graph attention networks (Velicković et al., 2017), which learn how to assign weights to a graph, but the graph adjacency structure is still fixed a priori through the masking in the attention mechanism.
@@
Previous work by Agrawal et al. (2019) studied backpropagation techniques for the differentiation of convex optimization problems. Our paper extends this work to nonlinear equations on graphs involving graph Laplacians.
@@
Deep learning methods have increasingly leveraged graph-based approaches or drawn inspiration from them. Two recent works propagate labels on a graph to generate pseudolabels in a semi-supervised setting (Iscen et al. (2019); Sellars et al. (2021)). However, the label propagation process is disconnected from the neural network and not directly integrated into the backpropagation or learning process. Contrastive learning (CL) uses pair-wise similarity, akin to a complete similarity graph, to facilitate the learning process (Chen et al. (2020b, a); Khosla et al. (2020); Wang et al. (2023)). Wang and Osher (2021) utilize graph learning as a classifier within a neural network, but they approximate the backpropagation gradients instead of computing them exactly.
@@
Figure 1: Comparison between the classic structure and our proposed graph learning layer (GLL) neural network structures. Both structures consist of a feature encoder and a classifier, but we replace the multilayer perceptron (MLP) classifier, consisting of one or more linear and activation layers and a final softmax classifier, with our data-dependent GLL with no trainable parameters. Another difference is that we subsample a “base” image set from the training data, which serves as the labeled nodes in graph learning classification.
@@ 1.2 Our Contributions
The most significant difference between our work and any of the preceding works is the derivation and implementation of the exact backpropagation equations through the graph learning layer, including the similarity graph construction and the graph-based classifier. We also give a sparse, efficient implementation of these equations. This is a significant extension of Enwright et al. (2023), where the authors derive the backpropagation gradients for Laplace learning in the special case of a fully connected graph. In this work, we derive the gradients for any $k$-nearest neighbors-type graph construction and any nonlinear elliptic graph learning equation. In addition, we conduct detailed experiments on a variety of benchmark datasets and neural network architectures. Our analysis includes visualizing clustering effects in the latent space under different GLL hyperparameters, demonstrating improved training dynamics and generalization, and showcasing improved robustness to adversarial attacks. In many experiments, GLL outperforms the standard projection head and softmax-based classifier.
The remainder of the paper is organized as follows. Section 2 details the necessary background on Laplace Learning, the seminal graph-based learning algorithm. Section 3 describes our novel mathematical contributions powering our novel GLL. Namely, we derive the gradient of the loss with respect to the similarity graph weight matrix and the associated $k$-Nearest Neighbor graph construction, which allows us to integrate graph learning into a neural network training context. We also demonstrate the effect this novel GLL has on the learned neural network embeddings. Section 4 contains illustrative experimental results showcasing the viability of the GLL on classification tasks with benchmark datasets, demonstrating strong classification accuracy as well as improved generalization and robustness to adversarial attacks compared to the softmax classifier.
@@ 2 Background
In this section, we give some background on construction similarity graphs, calculus on graphs, and graph learning using various types of graph Laplace equations.
@@ 2.1 Similarity Graph Construction
In this paper, we use an approximate $k$-nearest neighbor search using the Annoy python package [^2] https://github.com/spotify/annoy. The weights $w_{ij}$ are computed between the feature vectors of neighboring nodes $x_{i}$ and $x_{j}$ using the radial basis function
$w_{ij}=\exp\left(\frac{-4\|x_{i}-x_{j}\|^{2}}{\varepsilon_{k}(x_{i})\varepsilon_{k}(x_{j})}\right).$ (1)
The bandwidth parameter $\varepsilon_{k}(x_{i})$ is the distance from $x_{i}$ to its $k$th nearest neighbor, which allows the weights to self-tune appropriately to areas of density and sparsity. If desired for applications, $\varepsilon_{k}(x_{i})$ may be set to a constant value for all $i$. Because the k-nearest neighbor search is not necessarily symmetric, we symmetrize the weight matrix by replacing $w_{ij}$ with $\max(w_{ij},w_{ji})$.
@@ 2.2 Calculus on graphs
Let ${\mathcal{X}}$ be the vertex set of a graph with $n\in\mathbb{N}$ vertices and nonnegative edge weights $w_{xy}\geq 0$, for $x,y\in{\mathcal{X}}$. Let $W=(w_{xy})_{xy}$ denote the $n\times n$ weight matrix for the graph. We always assume the weight matrix $W$ is symmetric, so $w_{xy}=w_{yx}$.
@@
is an example of a vector field over the graph. For $\mathbf{u},\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$ we define an inner product
@@
The graph divergence is as an operator taking vector fields to functions in ${\ell^{2}({\mathcal{X}})}$, and is defined as the adjoint of the gradient. Here, we define the divergence for any function $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$; that is, it may not be a vector field. Hence, for $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$, the graph divergence $\operatorname{div}\mathbf{v}\in{\ell^{2}({\mathcal{X}})}$ is defined so that the identity
$\left\langle\nabla u,\mathbf{v}\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle u,\operatorname{div}\mathbf{v}\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}$ (6)
@@
which is the Dirichlet energy on the graph. For a connected graph with $k=1$, the kernel of the graph Laplacian $\Delta$ is $1$-dimensional and spanned by the vector of all ones. In general, the multiplicity of the eigenvalue 00 indicates the number of connected components in the graph (Von Luxburg (2007)).

### v2

@@ Abstract
Standard deep learning architectures used for classification generate label predictions with a projection head and softmax activation function. Although successful, these methods fail to leverage any relational information between samples for generating label predictions. In recent works, graph-based learning techniques, namely Laplace learning (zhu2003semi) or label propagation, have been heuristically combined with neural networks for both supervised and semi-supervised learning (SSL) tasks. However, prior works approximate the gradient of the loss function with respect to the graph learning algorithm or decouple the processes; end-to-end integration with neural networks is not achieved. In this work, we derive backpropagation equations, via the adjoint method, for the inclusion of a general family of graph learning layers into a neural network. The resulting method, distinct from graph neural networks, allows us to precisely integrate similarity graph construction and graph Laplacian-based label propagation into a neural network layer, replacing a projection head and softmax activation function for general classification tasks. Our experimental results demonstrate smooth label transitions across data, improved generalization and robustness to adversarial attacks, and improved training dynamics compared to a standard softmax-based approach [^1] Our code is available at https://github.com/jwcalder/GraphLearningLayer.
@@ 1 Introduction
Graph-based learning is a powerful class of transductive machine learning techniques that leverage graph structures in data for tasks such as semi-supervised learning, data visualization, and dimension reduction. Beginning with the seminal work by zhu2003semi on semi-supervised learning, graph-based learning techniques have since proliferated (belkin2004semi; belkin2004regularization; bengio2006label; zhou2011semi; wang2013dynamic; ham2005semisupervised; lee2013graph; calder2020poisson; calder2020properly; shi2017weighted) and found applications in many problem areas. In some settings, graphs are intrinsically present in a data set: links between servers in a network or between users on social media platforms are common in network science, while other problems may represent data as graphs themselves, such as molecules in drug discovery. In general, a similarity graph can be constructed over a data set by comparing pairs of data points using any available notion of similarity. In this paper, we focus on graph-based learning using similarity graphs; that is, we do not assume the graph is given as part of the metadata in the problem. Thus, our work is orthogonal and complementary to much of the graph neural network literature, where the graph is a priori fixed.
@@
Similarity graphs and corresponding graph techniques have been extensively studied theoretically using variational analysis, random walks, harmonic functions, and partial differential equations (see e.g., calder_consistency_2019; calder2022improved; calder2020properly; calder2018game; slepcev2019analysis; dunlop2020large; garcia2020error; hein2007graph; hein2005graphs; calder2022hamilton; calder2023rates). Graph learning also contrasts with modern deep learning approaches in that it is a data-dependent model without any trainable parameters. Previous works (chapman2023novel; chen2023batch; chen2023graphigarss) show that graph learning outperforms many popular classifiers like support vector machine (cortes1995support), random forest (ho1995random), and neural networks on classification tasks, especially at very low label rates.
@@
Recently, many works have focused on combining graph-based learning with deep learning. In such works, unsupervised or semi-supervised neural networks learn an embedding of the data into a feature space and use the similarity between features to construct the graph. Examples of such feature extraction networks include variational autoencoders (VAEs) (kingma2013auto; calder2020poisson; calder2022hamilton), convolutional neural networks (CNNs) (lecun1989backpropagation; enwright2023deep), convolutional VAEs (pu2016variational; miller2022graph; chapman2023novel), and contrastive learning (chen2020simple; brown2023utilizing; brown2023material). Some works have used graph learning to generate pseudolabels to train a deep neural network in low label rate settings (sellars2021laplacenet; iscen2019label). Of particular relevance to our work is a recent approach that replaces the softmax activation layer with graph learning to improve generalization and adversarial robustness (wang2021graph). However, wang2021graph do not compute the true gradients with respect to the loss, due to difficulties in backpropagating through graph learning, and instead replace them with a heuristic. Our work was partially motivated by the issues raised by wang2021graph.
@@
A significant gap exists in the literature as no work fully integrates graph learning into the gradient-based learning framework of neural networks. In this paper, we leverage the adjoint graph Laplace equation to derive the exact backpropagation gradients for similarity graph construction and a general family of graph learning methods, and show how to implement them efficiently. This significantly expands upon the previous work by enwright2023deep, where the backpropagation equations were only derived for Laplace learning on a fully connected graph. The backpropagation equations allow us to design a neural network layer that replaces the standard projection head and softmax classifier with a graph learning classifier (see Figure 2 for a comparison). We call this novel layer the Graph Learning Layer (GLL). A GLL (Figure 1) can be combined with any neural network architecture and enables end-to-end training in a gradient-based optimization framework. Through a variety of experiments, we show that the GLL learns the intrinsic geometry of data, improves the training dynamics and test accuracy of both shallow and deep networks across levels of supervision, and is significantly more robust to adversarial attacks compared to a projection head and softmax classifier.
@@ 1.1 Related Work
Graph-based semi-supervised learning (SSL) methods construct a graph that allows inferences of unlabeled nodes using labeled nodes and the graph topology. The seminal work of Laplace learning (zhu2003semi), also known as label propagation, led to a proliferation of related techniques (belkin2004semi; belkin2004regularization; bengio2006label; zhou2011semi; wang2013dynamic; ham2005semisupervised; lee2013graph; garcia2014multiclass; calder2020properly; shi2017weighted). Laplace learning computes a graph harmonic function to extend the given labels to the remainder of the graph, essentially finding the smoothest function that agrees with the given labels. When the label rate is low (close to one label per class), Laplace learning degenerates and suffers from a “spiking” phenomenon around labeled points, while predictions are roughly constant (calder2020poisson; el2016asymptotic; nadler2009infiniteunlabelled). Several methods have been proposed for these very low label rate regimes, including reweighted Laplace learning (shi2017weighted; calder2020properly; miller2023poisson), $p$-Laplace methods (zhou2011semi; el2016asymptotic; flores2022analysis; slepcev2019analysis), and Poisson learning (calder2020poisson). Analysis of the degeneracy of Laplace learning from a random walk and variational perspective can be found in (calder2020poisson; calder2023rates). Some recent work has also focused on graph total variation problems using graph cuts (bertozzi2016diffuse; merkurjev2017modified; merkurjev2018semi).
@@
Another related, but distinct, area of work in machine learning on graphs is graph neural networks, or GNNs (kipf2016semi; defferrard2016convolutional; wu2020comprehensive). The GNN literature considers how to incorporate a priori graph information relevant to the task, such as co-authorship in a citation graph, within a neural network architecture for various learning tasks, such as learning node embeddings or semi-supervised learning. The setting and problem space is quite different from ours. In particular, we do not assume the data has an inherent graph structure. Moreover, GNNs contain trainable weight matrices, while the graph-based learning methods we consider here have no trainable parameters. There are some works, such as graph attention networks (velivckovic2017graph), which learn how to assign weights to a graph, but the graph adjacency structure is still fixed a priori through masking in the attention mechanism.
@@
Previous work by agrawal2019differentiable studied backpropagation techniques for the differentiation of convex optimization problems. Our paper extends this work to nonlinear equations on graphs involving graph Laplacians.
@@
Deep learning methods have increasingly leveraged graph-based approaches or drawn inspiration from them. chen2024survey provides a broad overview of applications and modalities of graph structures and neural networks in computer vision. Two recent works propagate labels on a graph to generate pseudolabels in a semi-supervised setting (iscen2019label; sellars2021laplacenet). Other works have applied transformers directly to a generated similarity graph for node classification (wu2022nodeformer, han2022vision, zheng2022graph). However, in all of these works, the graph construction and label propagation process is disconnected from the neural network and is not directly integrated into the backpropagation or learning process. Contrastive learning (CL) uses pair-wise similarity, akin to a complete similarity graph, to facilitate the learning process (chen2020big; chen2020simple; khosla2020supervised; wang2023message). wang2021graph utilizes graph learning as a classifier within a neural network, but they approximate the backpropagation gradients with a linear layer instead of computing them exactly.
(a) Graph Learning Layer
(b) Standard MLP projection head
Figure 1: Visualization of the graph learning layer (GLL) within a neural network pipeline compared to a standard multilayer perceptron (MLP) projection head. For the GLL, a combination of labeled and unlabeled input images are batched and normalized into data for the feature extractor. The feature network encodes the images in feature space. The GLL is a combination of two steps; it generates a similarity graph from the encoded data and then propagates the labels across the graph, giving predictions on the unlabeled data. These predictions are used in the desired loss function and gradients can flow through the GLL to update the feature extractor. By contrast, the MLP projection head pipeline does not involve any initially labeled nodes and the classification method ignores relational information in the encodings. The + indicates learnable parameters in the network; note GLL has none.
@@
Figure 2: Comparison between the classic structure and our proposed graph learning layer (GLL) neural network structures for image classification. Both structures consist of a feature encoder and a classifier, but we replace the multilayer perceptron (MLP) classifier - consisting of one or more linear and activation layers and a final softmax classifier - with our data-dependent GLL with no trainable parameters. We reserve a subsample of the training data - denoted base samples - to use as the labeled nodes for label propagation. The samples that are not from the base set are use to compute the loss.
@@ 1.2 Our Contributions
The most significant difference between our work and any of the preceding works is the derivation and implementation of the exact backpropagation equations through the Graph Learning Layer, including the similarity graph construction and the graph-based classifier. We also give a sparse, efficient implementation of these equations. This is a significant extension of enwright2023deep, where the authors derive the backpropagation gradients for Laplace learning in the special case of a fully connected graph. In this work, we derive the gradients for any $k$-nearest neighbors-type graph construction and any nonlinear elliptic graph learning equation. In addition, we conduct detailed experiments on a variety of benchmark datasets and neural network architectures. Our analysis includes visualizing clustering effects in the latent space under different GLL hyperparameters, demonstrating improved training dynamics and generalization, and showcasing improved robustness to adversarial attacks. In the vast majority experiments - especially at low label rates - a GLL outperforms the standard projection head and softmax-based classifier.
The remainder of the paper is organized as follows. Section 2 details the necessary background on Laplace Learning, the seminal graph-based learning algorithm. Section 3 describes our novel mathematical contributions powering the GLL. Namely, we derive the gradient of the loss with respect to the similarity graph weight matrix and the associated $k$-Nearest Neighbor graph construction, which allows us to integrate graph learning into a neural network training context. We also demonstrate the effect this novel GLL has on the learned neural network embeddings. Section 4 contains illustrative experiments showcasing the benefits of the GLL, demonstrating improved generalization across network architectures and levels of supervision, and superior robustness to adversarial attacks compared to the softmax classifier.
@@ 2 Background
In this section, we give some background on the construction similarity graphs, calculus on graphs, and graph learning using various types of graph Laplace equations.
@@ 2.1 Similarity Graph Construction
Let $\mathcal{X}$} be the vertex set of a graph with $n\in\mathbb{N}$ vertices and nonnegative edge weights $w_{xy}\geq 0$, for $x,y\in{\mathcal{X}}$. Let $W=(w_{xy})_{x,y\in{\mathcal{X}}}$ denote the $n\times n$ weight matrix for the graph. We always assume the weight matrix $W$ is symmetric, so $w_{xy}=w_{yx}$. We usually think of $W$ as a similarity matrix, in which case the entry $w_{xy}$ encodes how similar the nodes $x$ and $y$ are, with large positive values indicating a high degree of similarity, and $w_{xy}=0$ indicating the nodes are not similar. Given we have a feature vector in $\mathbb{R}^{d}$ for each node $x\in{\mathcal{X}}$ in the graph, a common way to construct a similarity weight matrix, which we adopt in the numerical experiments in this paper, is a $k$-nearest neighbor graph based on similarity of feature vectors. In this case, we will denote the vertex set of the graph by ${\mathcal{X}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}\in\mathbb{R}^{d}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is the feature vector for the $i^{\rm th}$ node in the graph. The weight $w_{xy}=w_{\mathbf{x}_{i}\mathbf{x}_{j}}=w_{ij}$ between nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is then given by the expression
$w_{ij}=\exp\left(\frac{-4\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{\varepsilon_{k}(\mathbf{x}_{i})\varepsilon_{k}(\mathbf{x}_{j})}\right).$ (1)
The parameter $k$ is the number of neighbors used in the graph construction, and we choose the bandwidth parameter $\varepsilon_{k}(\mathbf{x}_{i})$ to be the distance from $\mathbf{x}_{i}$ to its $k$th nearest neighbor in $\mathbb{R}^{d}$, which allows the weights to self-tune appropriately to areas of density and sparsity. If desired for applications, $\varepsilon_{k}(\mathbf{x}_{i})$ may be set to a constant value for all $\mathbf{x}_{i}$. To produce a sparse graph, we set the weight $w_{ij}=0$ if $\mathbf{x}_{i}$ is not among the $k$ nearest neighbors of $\mathbf{x}_{j}$, and $\mathbf{x}_{j}$ is not among the $k$ nearest neighbors of $\mathbf{x}_{i}$. While searching for the exact $k$ nearest neighbors in high dimensional data sets is a computational hard problem, there are many fast approximate nearest neighbor algorithms that can be used in practice. [^2] We use the Annoy python package: https://github.com/spotify/annoy While similarity matrices are often used in practice, we note that the theory in this section, and much of Section 3, does not place any restrictions on $W$ aside from symmetry $w_{xy}=w_{yx}$.
@@ 2.2 Calculus on graphs
is an example of a vector field over the graph. Note that $\nabla u$ is a vector field, so it is defined over the edges in the graph, i.e., pairs $(x,y)$ of vertices, while $u$ is a function of a single vertex $x$ on the graph. For $\mathbf{u},\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$ we define an inner product
@@
The graph divergence is an operator taking vector fields to functions in $\ell^{2}({\mathcal{X}})$}, and is defined as the adjoint of the gradient. Here, we define the divergence for any function $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$; that is, it may not be a vector field. Hence, for $\mathbf{v}\in{\ell^{2}({\mathcal{X}}^{2})}$, the graph divergence $\operatorname{div}\mathbf{v}\in{\ell^{2}({\mathcal{X}})}$ is defined so that the identity
$\left\langle\nabla u,\mathbf{v}\right\rangle_{{\ell^{2}({\mathcal{X}}^{2})}}=\left\langle u,\operatorname{div}\mathbf{v}\right\rangle_{{\ell^{2}({\mathcal{X}})}}$ (6)
@@
which is the Dirichlet energy on the graph. For a connected graph with $k=1$, the kernel of the graph Laplacian $\Delta$ is $1$-dimensional and spanned by the vector of all ones. In general, the multiplicity of the eigenvalue 0 indicates the number of connected components in the graph (von2007tutorial).